# Revisiting Semi-Supervised Learning in the Era of Foundation Models

### v1

@@ Abstract
Semi-supervised learning (SSL) leverages abundant unlabeled data alongside limited labeled data to enhance learning. As vision foundation models (VFMs) increasingly serve as the backbone of vision applications, it remains unclear how SSL interacts with these pre-trained models. To address this gap, we develop new SSL benchmark datasets where frozen VFMs underperform and systematically evaluate representative SSL methods. We make a surprising observation: parameter-efficient fine-tuning (PEFT) using only labeled data often matches SSL performance, even without leveraging unlabeled data. This motivates us to revisit self-training, a conceptually simple SSL baseline, where we use the supervised PEFT model to pseudo-label unlabeled data for further training. To overcome the notorious issue of noisy pseudo-labels, we propose ensembling multiple PEFT approaches and VFM backbones to produce more robust pseudo-labels. Empirical results validate the effectiveness of this simple yet powerful approach, providing actionable insights into SSL with VFMs and paving the way for more scalable and practical semi-supervised learning in the era of foundation models. [^1] Our code is available at https://github.com/OSU-MLB/SSL-Foundation-Models.
semi-supervised learning, self-training, vision foundation models, foundation models, benchmark, parameter-efficient fine-tuning
@@ 1 Introduction
The quality of machine learning models is often closely tied to the amount of labeled data available, but annotation can be costly or labor-intensive. As a result, semi-supervised learning (SSL), which leverages abundant unlabeled data alongside limited labeled data, has emerged as a promising paradigm for developing ML models without the need for extensive labeling (Zhu, 2008; Chapelle et al., 2006; Van Engelen & Hoos, 2020; Yang et al., 2022). Over the past few decades, numerous SSL algorithms have been developed to advance this field. Exemplar methods from the deep learning era include Mean Teacher (Tarvainen & Valpola, 2018), MixMatch (Berthelot et al., 2019), and FixMatch (Sohn et al., 2020), which dynamically impose objective functions on unlabeled data based on distillation and consistency, thereby enhancing learning performance. It is worth noting that many of these methods were originally designed to train neural networks from scratch.
Meanwhile, pre-training on external labeled or unlabeled data has become the de facto standard across many machine learning application domains (Gui et al., 2024). For example, in computer vision, many recent algorithms are built upon vision foundation models (VFMs) such as CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2024). These models were pre-trained on massive datasets—–millions, if not billions, of data points. As a result, they have demonstrated remarkable generalizability across a wide range of tasks, often requiring only slight fine-tuning or, in some cases, functioning effectively as a frozen backbone (Zhang et al., 2024).
Figure 1: Illustration of our proposed SSL pipeline. To effectively leverage abundant unlabeled data alongside scarce labeled data in the era of VFMs, our approach follows four phases: (1) Supervised Parameter-Efficient Fine-Tuning, where we harness labeled data by fine-tuning pre-trained VFMs using various PEFT algorithms; (2) Pseudo-Label Generation, where we exploit fine-tuned VFMs’ generalization ability to generate pseudo-labels for unlabeled data; (3) Pseudo-Label Ensemble, where we enhance robustness by aggregating pseudo-labels from multiple fine-tuned VFMs; and (4) Self-Training, where we consolidate all knowledge into one model.
@@
To this end, we introduce new SSL benchmark datasets based on the Visual Task Adaptation Benchmark (VTAB) (Zhai et al., 2020), a diverse suite of classification tasks designed to evaluate visual representations. Our focus is on tasks where frozen VFMs underperform, indicating the need for further fine-tuning, and where SSL could offer a beneficial solution. We then systematically evaluate three recent representative SSL methods—FixMatch (Sohn et al., 2020), FlexMatch (Zhang et al., 2022), SoftMatch (Chen et al., 2023) and FineSSL (Gan & Wei, 2024). Hyperparameters are carefully selected using techniques proposed for unsupervised domain adaptation (Ericsson et al., 2023), ensuring that the data leakage issue discussed in (Oliver et al., 2018) is avoided.
Our empirical results show that SSL does not offer a significant advantage over fully fine-tuning VFMs with limited labeled data. In some cases, accuracy even decreases. We hypothesize that allowing SSL to update all parameters in VFMs may inadvertently reduce their built-in generalizability. Indeed, in practical applications of VFMs, parameter-efficient fine-tuning (PEFT)—which updates only a small fraction of parameters or introduces a lightweight learnable module while keeping the VFM largely frozen—often outperforms full fine-tuning, particularly when downstream labeled data is limited (Mai et al., 2024; Xin et al., 2024). When we integrate PEFT into SSL, performance improves significantly. However, even when PEFT alone is used to fine-tune VFMs, the performance gains often match or exceed those by SSL, all without requiring unlabeled data.
These findings convey two key messages. First, new SSL algorithms specifically designed for VFMs are needed. Second, the fact that supervised PEFT matches SSL accuracy suggests it can already generate high-quality pseudo-labels for unlabeled data. Appropriately leveraging these pseudo-labels could further improve model performance.
Building on these insights, we revisit self-training, a conceptually simple SSL baseline that uses the current model to pseudo-label unlabeled data for further training (Lee et al., 2013; Xie et al., 2020; Chen et al., 2020; Yalniz et al., 2019). A key challenge in self-training is ensuring the reliability of pseudo-labels—–incorrect labels can reinforce model biases and degrade performance. Traditional approaches rely on confidence thresholds or consistency constraints to filter pseudo-labels (Sohn et al., 2020; Nguyen & Yang, 2023), but these methods are often heuristic and suboptimal, limiting self-training performance compared to more advanced SSL methods.
To address this, we propose an alternative strategy that leverages the complementary behaviors of VFMs and PEFT. Different VFMs and PEFT algorithms often produce diverse predictions, even when their overall accuracies are  similar (Mai et al., 2024; Tong et al., 2024b).
This inherent diversity enables us to harness ensembling techniques (Zhou, 2012). By aggregating predictions from multiple widely used VFM backbones fine-tuned with different PEFT methods—and explicitly compensating for their varied confidence distributions—we obtain significantly more robust pseudo-labels, reducing the need for explicit pseudo-label selection.
As a result, our ensemble-based approach yields a more reliable yet simpler self-training pipeline (Figure 1) that effectively leverages labeled and unlabeled data as well as the strengths of VFMs and PEFT, leading to substantial improvements over existing SSL methods.
We extensively validate our approach on new benchmark datasets, where it outperforms existing SSL methods across most tasks, establishing it as a simple, effective, and competitive SSL baseline in the era of foundation models. Further analyses provide key insights for future improvements.
@@
Our key contributions are as follows:
  • We introduce new SSL benchmark datasets tailored for studying SSL in the context of foundation models.
  • We conduct a systematic evaluation of representative SSL methods, offering valuable insights for future research.
  • We propose a simple yet highly effective self-training-based SSL method leveraging the key properties of VFMs and PEFT, establishing a strong baseline for future work.
Remark. Self-training (Amini et al., 2025), pseudo-labeling (Lee et al., 2013), and ensembling (Dong et al., 2020) have been extensively studied in the ML literature, spanning both the pre-deep learning and deep learning eras. Our goal is not to compete with existing methods but to establish a simple yet effective semi-supervised learning baseline that harnesses their strengths while incorporating the unique properties of foundation models. By maintaining a clean and accessible approach, we aim to facilitate future research. While we acknowledge that integrating more sophisticated techniques could further enhance performance, we deliberately focus on simplicity.
A key step in ensembling is obtaining multiple base learners that are both diverse and equally performant. Our contribution lies in leveraging the complementary behaviors of multiple foundation models and PEFT methods—an approach tailored specifically for the foundation model era.
2 Related Works
Semi-Supervised Learning. In the past few years, numerous SSL algorithms have been developed, many of which focus on generating and selecting reliable pseudo-labels during the training process (Sohn et al., 2020; Berthelot et al., 2019, 2020). For example, FixMatch (Sohn et al., 2020) relies on a fixed threshold to select pseudo-labels, while more recent methods dynamically adjust this threshold as training proceeds. FlexMatch (Zhang et al., 2022) introduces class-specific thresholds to accommodate varying degrees of difficulty, SoftMatch (Chen et al., 2023) applies a soft confidence threshold to balance label quality and quantity. However, most existing approaches were originally designed to train models from scratch, leaving open questions about their compatibility with VFMs and the adjustment required for optimal performance. Another prominent direction in SSL is self-training (Xie et al., 2020; Lee et al., 2013), typically comprising three steps: (1) train a teacher model on labeled data, (2) generate pseudo-labels on unlabeled data using the teacher, and (3) train a student model on both labeled and pseudo data. A key challenge lies in ensuring the reliability of pseudo-labels. Conventional solutions employ unsupervised pre-training (Chen et al., 2020), confidence thresholds or consistency constraints (Sohn et al., 2020; Nguyen & Yang, 2023) to improve the pseudo-label quality. In contrast, we aim to leverage the power of VFMs to produce more robust pseudo-labels and turn self-training into a simple yet powerful baseline in the VFM era.
Vision Foundation Models. Vision Transformers (ViT) (Dosovitskiy et al., 2021) pre-trained on massive amounts of data, has become indispensable to modern AI development (Awais et al., 2025). These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks. For example, CLIP-ViT (Radford et al., 2021) with millions of image-text pairs show an unprecedented zero-shot capability, robustness to distribution shifts and serve as the encoders for various power generative models (Rombach et al., 2022; Liu et al., 2024)). Meanwhile, DINOv2 (Oquab et al., 2024) pre-trained with self-supervised objectives on extensive sets of well-curated images effectively captures fine-grained localization features. Despite this growing prominence of VFMs, limited work has explored how to integrate them with SSL. Although a recent study marks an initial attempt (Gan & Wei, 2024), a comprehensive study of how to effectively leverage VMFs in SSL is still lacking.
@@ 3.1 Problem Definition
We consider a $C$-class classification problem in SSL with a large unlabeled dataset $\mathcal{D}^{u}\coloneqq\{{\bm{x}}^{u}_{i}\}_{i=1}^{N^{u}}$ and a much smaller labeled dataset $\mathcal{D}^{l}\coloneqq\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{N^{l}}$, where ${\bm{x}}^{l}_{i}$ and ${\bm{x}}^{u}_{i}$ are training samples and $y^{l}_{i}$ is the ground-truth class label. $N^{u}$ and $N^{l}$ denote the unlabeled and labeled dataset size with $N^{u}\gg N^{l}$. SSL aims to learn a model $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ using $\mathcal{D}^{u}$ and $\mathcal{D}^{l}$. Unlike conventional SSL that initializes $\bm{\theta}$ randomly, our framework starts with a VFM (e.g., CLIP or DINOv2) and fine-tunes it for downstream tasks.
3.2 A Diverse and Challenging SSL Benchmark
Despite recent advancements in SSL, most studies (Gan & Wei, 2024; Chen et al., 2023; Zhang et al., 2022) still evaluate performance on classic datasets such as CIFAR-10/100 (Krizhevsky & Hinton, 2009), STL-10 (Coates et al., 2011), and Food101 (Bossard et al., 2014). These benchmarks were suitable when models were randomly initialized, but with the advent of VFMs, they have become far less challenging. As shown in Table 1, linear probing on top of frozen VFM backbones already achieves remarkable accuracy on these benchmarks, even with fairly limited labeled data. Moreover, these datasets primarily focus on natural images, providing a limited perspective on real-world scenarios where VFMs are deployed.
Table 2: Summary of our benchmark. Rem. Sen.: Remote sensing; Recog.: Recognition; $N^{l}/N^{u}$: # labeled/unlabeled training data.
Dataset Task Domain $N^{u}+N^{l}$ Classes $N^{l}/$Class

### v2

@@ Abstract
Semi-supervised learning (SSL) enhances model performance by leveraging abundant unlabeled data alongside limited labeled data. As vision foundation models (VFMs) become central to modern vision applications, this paper revisits SSL in the context of these powerful pre-trained models. We conduct a systematic study on tasks where frozen VFMs underperform and reveal several key insights when fine-tuning them. First, parameter-efficient fine-tuning (PEFT) using only labeled data often surpasses traditional SSL methods—even without access to unlabeled data. Second, pseudo-labels generated by PEFT models offer valuable supervisory signals for unlabeled data, and different PEFT techniques yield complementary pseudo-labels. These findings motivate a simple yet effective SSL baseline for the VFM era: ensemble pseudo-labeling across diverse PEFT methods and VFM backbones. Extensive experiments validate the effectiveness of this approach, offering actionable insights into SSL with VFMs and paving the way for more scalable and robust semi-supervised learning in the foundation model era.
footnotetext: *Equal contribution.footnotetext: Our code is available at https://github.com/OSU-MLB/SSL-Foundation-Models.
@@ 1 Introduction
The quality of machine learning (ML) models is often closely tied to the amount of labeled data available, but annotation can be costly or labor-intensive. Semi-supervised learning (SSL), which leverages abundant unlabeled data alongside limited labeled data, has thus emerged as a promising paradigm for developing ML models without the need for extensive labeling zhu2008semi ; chapelle2006semi ; van2020survey ; yang2022survey . Over the past few decades, numerous SSL algorithms have been developed to advance this field. Exemplar methods from the deep learning era include Mean Teacher tarvainen2018meanteachersbetterrole , MixMatch berthelot2019mixmatch , and FixMatch sohn2020fixmatch , which dynamically impose objective functions on unlabeled data based on distillation and consistency, thereby enhancing learning performance. It is worth noting that many of these methods were originally designed to train neural networks “from scratch.”
Figure 1: Left: The Venn diagram of the top 20% highest-confidence predictions from various VFMs and PEFT under DTD 3-shot, illustrating their intrinsic property of producing a diverse range of high-confidence predictions. Right: Ensembling these diverse pseudo-label predictions progressively boosts downstream performance for increasingly more ensembles (Self-Training → PETV-PET), highlighting the quality improvements from diversity. Results are averaged across 12 settings in our benchmark.
Recently, pre-training on external labeled or unlabeled data has become the de facto standard across many machine learning application domains gui2024survey . For example, in computer vision, many recent algorithms are built upon vision foundation models (VFMs) such as CLIP radford2021learningtransferablevisualmodels and DINOv2 oquab2024dinov2learningrobustvisual . These models were pre-trained on massive datasets—millions, if not billions, of data points. As a result, they have demonstrated remarkable generalizability across a wide range of tasks, often requiring only slight fine-tuning or, in some cases, functioning effectively as a frozen backbone zhang2024vision .
@@
Study design. To this end, we introduce new SSL benchmark datasets based on the Visual Task Adaptation Benchmark (VTAB) zhai2020largescalestudyrepresentationlearning , a diverse suite of classification tasks designed to evaluate visual representations. Our focus is on tasks where frozen VFMs underperform, indicating the need for further fine-tuning, and where SSL could offer a beneficial solution. We then systematically evaluate three representative SSL methods—FixMatch sohn2020fixmatch , FlexMatch zhang2021flexmatch and SoftMatch chen2023softmatch . Hyperparameters are carefully selected using techniques proposed for unsupervised domain adaptation ericsson2023betterpracticesdomainadaptation , ensuring that the data leakage issue discussed in oliver2018realistic is avoided.
Key insights. Our empirical results highlight two main findings. First, fine-tuning VFMs with representative SSL algorithms offers limited advantages over using only labeled data for fine-tuning. Second, parameter-efficient fine-tuning (PEFT) mai2024lessonslearnedunifyingempirical ; vpetlbench2024 —which updates a small subset of parameters or adds lightweight learnable modules while keeping the VFM largely frozen—consistently yields substantial performance gains, regardless of the learning paradigm.
These observations imply two important directions. First, there is a need for SSL methods tailored specifically for VFMs. Second, since PEFT models trained only on labeled data already match the performance of standard SSL approaches, effectively leveraging their predictions—i.e., pseudo-labels—offers a promising path to further improve performance in semi-supervised settings.
An SSL baseline in the VFM era. Building on these insights, we introduce a simple yet effective SSL approach that leverages VFMs as backbones. Our method is grounded in the principle of self-training—a straightforward SSL strategy in which the model generates pseudo-labels for unlabeled data to guide further training lee2013pseudo ; xie2020selftrainingnoisystudentimproves ; chen2020bigselfsupervisedmodelsstrong ; yalniz2019billionscalesemisupervisedlearningimage . While traditional self-training often struggles with low-quality pseudo-labels, we address this by exploiting two key properties of VFMs and PEFT methods: their strong initial performance and complementary behaviors.
Specifically, as observed in mai2024lessonslearnedunifyingempirical ; tong2024eyes , different VFM backbones and PEFT methods frequently produce diverse predictions—even when their overall accuracies are similar. As shown in Figure 1, this diversity motivates ensembling predictions from multiple VFM-PEFT pairs zhou2012ensemble . By explicitly compensating for their varied confidence distributions, we obtain significantly more robust pseudo-labels—without requiring explicit filtering sohn2020fixmatch ; nguyen2023boostingsemisupervisedlearningbridging . The result is a simpler yet more reliable self-training pipeline (Figure 2) that effectively leverages both labeled and unlabeled data, achieving substantial gains over existing SSL methods.
We extensively validate our approach, VFM-PEFT Ensemble Training (V-PET), on newly proposed benchmark datasets. V-PET consistently outperforms existing SSL methods across most tasks, including the recently proposed FineSSL gan2024erasingbiasfinetuningfoundation , which also builds on VFMs. These results establish PET as a simple, effective, and competitive SSL baseline for the foundation model era.
Figure 2: Illustration of V-PET. To effectively leverage abundant unlabeled data alongside scarce labeled data in the era of VFMs, our approach follows four phases: (a) Supervised Parameter-Efficient Fine-Tuning, where we harness labeled data by fine-tuning pre-trained VFMs using various PEFT algorithms; (b) Pseudo-Label Generation, where we exploit fine-tuned VFMs’ generalization ability to generate pseudo-labels for unlabeled data; (c) Pseudo-Label Ensemble, where we enhance robustness by aggregating pseudo-labels from multiple fine-tuned VFMs; and (d) Self-Training, where we consolidate all knowledge into one model.
Remark. Self-training amini2025self ; assran2022maskedsiamesenetworkslabelefficient , pseudo-labeling lee2013pseudo , and ensemble methods dong2020survey have been extensively studied in the ML literature, spanning both the pre-deep learning and deep learning eras. Our goal is not to compete with existing methods but to establish a simple yet effective semi-supervised learning baseline that harnesses their strengths while incorporating the unique properties of foundation models. We note that a key step in ensembling is obtaining multiple base learners that are diverse and equally performant. Our contribution lies in leveraging the complementary behaviors of multiple foundation models and PEFT methods—an approach tailored specifically for the foundation model era.
Consistent with representative SSL studies berthelot2019mixmatch ; sohn2020fixmatch ; zhang2021flexmatch , our primary focus is on classification. Nonetheless, we believe that our insights are transferable to various tasks and hope they will inspire future research on other downstream SSL tasks, such as segmentation and detection.
2 Related Works
Semi-Supervised Learning (SSL). In recent years, many SSL methods have centered on generating and selecting reliable pseudo-labels sohn2020fixmatch ; berthelot2019mixmatch ; berthelot2020remixmatch . FixMatch sohn2020fixmatch uses a fixed confidence threshold, while FlexMatch zhang2021flexmatch adopts class-specific thresholds and SoftMatch chen2023softmatch applies a soft threshold to balance label quality and quantity. However, these approaches were designed for training from scratch, leaving open questions about their compatibility with VFMs. Another key paradigm is self-training xie2020selftrainingnoisystudentimproves ; lee2013pseudo , which (1) trains a teacher on labeled data, (2) generates pseudo-labels for unlabeled samples, and (3) trains a student on both. Ensuring pseudo-label reliability typically involves unsupervised pre-training chen2020bigselfsupervisedmodelsstrong , confidence thresholds, or consistency constraints sohn2020fixmatch ; nguyen2023boostingsemisupervisedlearningbridging . We instead leverage VFMs to produce more robust pseudo-labels, questioning whether scratch-oriented SSL algorithms remain effective in the foundation-model era.
Transfer Learning & Self-Supervised Learning. Transfer learning zhuang2020comprehensivesurveytransferlearning —and specifically PEFT (or PETL) for efficiently adapting foundation models han2024parameterefficientfinetuninglargemodels —has long leveraged pretrained models to boost downstream tasks. In our approach, we fine-tune VFMs via PEFT on labeled data, then use the adapted models to generate pseudo-labels for unlabeled examples, thus bridging transfer learning with semi-supervised methods. A related paradigm, self-supervised learning, also exploits unlabeled data and often serves as a baseline in SSL comparisons chen2020bigselfsupervisedmodelsstrong ; assran2022maskedsiamesenetworkslabelefficient , but it assumes abundant unlabeled data—an assumption that may fail in low-resource settings.
Vision Foundation Models. Vision Transformers (ViT) dosovitskiy2021imageworth16x16words pre-trained on massive amounts of data, has become indispensable to modern AI development awais2025foundation . These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks. For example, CLIP-ViT radford2021learningtransferablevisualmodels trained with millions of image-text pairs shows an unprecedented zero-shot capability, robustness to distribution shifts, and serves as the encoders for various powerful generative models rombach2022high ; liu2024visual . Meanwhile, DINOv2 oquab2024dinov2learningrobustvisual pre-trained with self-supervised objectives on extensive sets of well-curated images effectively captures fine-grained localization features. Moreover, it is increasingly becoming a common sense that strategically fusing multiple VFMs can unlock synergistic gains mai2024lessonslearnedunifyingempirical ; tong2024eyes , leading to state-of-the-art performance across tasks from visual question answering to object detection.
Leveraging VFMs for Semi-Supervised Learning. With the growing prominence of VFMs, few recent studies have explored their usage in SSL. The most relevant to ours is FineSSL gan2024erasingbiasfinetuningfoundation , which specifically investigates the use of CLIP vision backbone in an SSL setting, accompanied by pseudo-labels refinement using a balanced margin softmax. The evaluation was limited to simple, small-scale datasets like CIFAR-10 krizhevsky2009learning , where frozen VFMs can already achieve high accuracy. A few other works constrained the study only to VFMs with zero-shot capabilities, such as the full CLIP model with both vision and text encoders menghini2024enhancingclipclipexploring ; zhang2024candidatepseudolabellearningenhancing . Our work, in contrast, extends the scope in both the diversity of VFMs and the evaluation benchmarks, establishing the first comprehensive study of SSL in the era of foundation models.
@@ 3.1 Problem Definition
We consider a $C$-class classification problem in SSL with a large unlabeled dataset $\mathcal{U}\coloneqq\{{\bm{x}}^{u}_{i}\}_{i=1}^{|\mathcal{U}|}$ and a much smaller labeled dataset $\mathcal{L}\coloneqq\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{|\mathcal{L}|}$, where ${\bm{x}}^{l}_{i}$ and ${\bm{x}}^{u}_{i}$ are training samples and $y^{l}_{i}$ is the ground-truth class label. $|\mathcal{U}|$ and $|\mathcal{L}|$ denote the unlabeled and labeled dataset size with $|\mathcal{U}|\gg|\mathcal{L}|$. SSL aims to learn a model $f_{\bm{\theta}}$ parameterized by $\bm{\theta}$ using $\mathcal{U}$ and $\mathcal{L}$. Unlike conventional SSL that initializes $\bm{\theta}$ randomly, our framework starts with a VFM (e.g., , CLIP or DINOv2) and fine-tunes it for downstream tasks.
3.2 A Comprehensive SSL Image Classification Benchmark
Despite recent advancements in SSL, most studies gan2024erasingbiasfinetuningfoundation ; chen2023softmatch ; zhang2021flexmatch continue to evaluate on classic datasets such as CIFAR-10/100 krizhevsky2009learning , STL-10 pmlr-v15-coates11a , and Food101 bossard2014food . However, these benchmarks exhibit two key limitations: Diminishing difficulty under VFMs. Following prior PEFT-on-VFM work jie2022convolutionalbypassesbettervision ; jia2022visual ; tu2023visual ; mai2024lessonslearnedunifyingempirical , we use linear probing as our primary performance indicator. As shown in Table 1, linear probing on frozen VFM backbones already delivers remarkable accuracy—even with limited labeled data. We also report results for various classification-head sizes in section B.4, and Narrow domain coverage. Because these benchmarks focus mainly on natural images, they offer only a narrow view of real-world VFM applications. To effectively evaluate SSL methods in the VFM era—focusing exclusively on semi-supervised image classification—we propose a new benchmark designed to capture the complexities of real-world applications across different domains and dataset sizes.
Dataset & Regime We choose six classification datasets from VTAB zhai2020largescalestudyrepresentationlearning , spanning remote sensing, medical imaging, synthetic data, and autonomous driving tasks. To test the robustness of SSL methods, we vary the number of labels per class and employ linear probing as our performance metric, selecting shot counts that keep each task sufficiently challenging for frozen VFMs. These settings impose substantial difficulty on frozen VFMs (Table 1), underscoring the importance of SSL for unlocking their full potential. A summary of the datasets appears in Table 2, with full details provided in Appendix A. Together, this benchmark offers a more diverse and comprehensive foundation-model-era evaluation of semi-supervised image classification.
3.3 Fair Hyperparameter Tuning for SSL
Hyperparameter tuning has long been a persistent challenge in SSL oliver2018realistic , remaining ambiguous and unstandardized throughout much of the existing literature. Due to the labeled data scarcity, a standard train-validation split is often infeasible. Tuning on a held-out labeled validation or even test set can lead to significant data leakage, resulting in over-optimistic results and unfair comparisons. To tackle this issue and complement our benchmark, we establish a more rigorous protocol for hyperparameter tuning in SSL.
Table 2: Summary of our benchmark. Rem. Sen.: Remote sensing; Recog.: Recognition; $|\mathcal{L}|$: # labeled training data; $|\mathcal{U}|$: # unlabeled training data.
Dataset Task Domain $|\mathcal{L}|+|\mathcal{U}|$ Classes $|\mathcal{L}|/$Class